# Machine Learning for Two-Sample Testing under Right-Censored Data: A Simulation Study

### v1

@@ 7 Conclusion
In this study, we explored how machine learning methods can be applied to solve two-sample problems with right-censored observations. We formulated the solution to the problem as a stacking of classical two-sample tests (in fact, using machine learning methods as a selection function). To train and test this method, we prepared a synthetic dataset using Monte Carlo simulations.
Despite the lack of the limit distribution for calculating the $p$-value (this issue can be resolved using the Monte Carlo simulation), the results indicate that the proposed methods enhance the accuracy of statistical inferences in two-sample testing, whether under right-censored data or not. Testing this hypothesis is an important task in many areas, such as medicine, sociology, engineering, and many others.
To be able to reproduce the results and use the proposed methods, we published the original synthetic dataset, including the source code for modeling (Hugging Face), and the trained models, including the source code for running (GitHub).
@@
The supplementary materials contain: (1) Table with absolute values of simulated test power; (2) Table with percentage points of the proposed methods including an averaged error of the $p$-value.
@@
The dataset simulated during and/or analyzed during the current study is available in the Hugging Face repository (pfilonenko/ML_for_TwoSampleTesting/), DOI: 10.57967/hf/2978.
The source code for the dataset simulation is available in the GitHub repository (pfilonenko/ML_for_TwoSampleTesting/dataset/simulation/).
The source code for the proposed methods and trained models are available in the GitHub repository (pfilonenko/ML_for_TwoSampleTesting).

### v2

@@ 6 Conclusion
In this paper, we demonstrated that the application of machine learning methods to solve the two-sample problem with right-censored observations can significantly enhance the accuracy of statistical inferences. To achieve this, we formulated the solution to the problem as a stacking of classical two-sample tests using machine learning methods as a selection function.
We investigated the statistical power and behavior of the null distribution of the proposed methods. To conduct our numerical experiments, we prepared a synthetic dataset using Monte Carlo simulations. Despite the absence of the limit distribution necessary for the $p$-value computation (this issue can be resolved using Monte Carlo simulation), the results indicate that the proposed methods improve the accuracy of statistical inferences in two-sample testing, whether under right-censored data or not.
The results obtained can find applications in many fields of science (engineering, chemistry, physics, biology, medicine, etc.) where comparison of two samples with right-censored (incomplete) observations is required. The use of the proposed methods will enhance the accuracy of decision-making in these applied areas.
To be able to reproduce the results or apply the proposed methods, we have published the original synthetic dataset, including the source code (Hugging Face), and trained models, including example scripts (GitHub).
@@
The supplementary materials contain: (1) Table with absolute power values; (2) Table with percentage points for the proposed methods.
@@
The dataset simulated during and analyzed during the current study is available in the Hugging Face repository, DOI: 10.57967/hf/2978.
The source code of the proposed methods is available in the GitHub repository:
https://github.com/pfilonenko/ML_for_TwoSampleTesting.